OpenStudy (anonymous):

Curtis is looking over the stat book from his basketball game yesterday. He noticed that he made 18 shots during the course of the game for a total of 41 points. If Curtis took both two-point shots and three-point shots, how many of each shot did he make?

OpenStudy (anonymous):

Justin bieber

OpenStudy (anonymous):

@Agent_Sniffles help me please?

OpenStudy (bahrom7893):

x+y=18 2x+3y=41

OpenStudy (bahrom7893):

Solve the system, call me when you're done and I'll take a look and see if it's correct.

OpenStudy (anonymous):

@bahrom7893 x=26 and y=5

OpenStudy (bahrom7893):

So according to you 26+5=18? Try again.

OpenStudy (anonymous):

@bahrom7893 x=13 and y=5 =D

OpenStudy (anonymous):

@bahrom7893 how would i put this in a sentence?